OpenStudy (anonymous):

What is the least positive integer k for which 15k is the cube of a number?

OpenStudy (asnaseer):

first write 15 as a product of primes

OpenStudy (anonymous):

3 times 5

OpenStudy (asnaseer):

good. now can you see what 'k' must be in order to get 15k to be a cube?

OpenStudy (asnaseer):

you have: 15k = 3 * 5 * k

OpenStudy (anonymous):

and then ...

OpenStudy (asnaseer):

ok - first just concentrate on the '3'. what must you multiply '3' by in order to get a perfect cube?

OpenStudy (anonymous):

3 times 3 times 3

OpenStudy (asnaseer):

perfect! can you what to do now?

OpenStudy (asnaseer):

*you see

OpenStudy (anonymous):

225?

OpenStudy (anonymous):

5 timse 5 times 5?

OpenStudy (asnaseer):

5^3 = 125

OpenStudy (asnaseer):

so, to make '3' into a cube (smallest cube) - you need to multiply it by 3^2 similarly, to make '5' into a cube - you need to multiply it by 5^2

OpenStudy (anonymous):

OH I GET IT NOW

OpenStudy (anonymous):

u need to cube 15, but since theires oalready 1 15, you need two more, or 15 squared which is 125

OpenStudy (asnaseer):

you must remember to first express the number as a product of its prime factors

OpenStudy (asnaseer):

e.g. what is the smallest cube for 45k ?

OpenStudy (asnaseer):

it is not 45^3

OpenStudy (anonymous):

3375?

OpenStudy (asnaseer):

45 = 3^2 * 5 so smallest cube would be achieved by multiplying 45 by 3 * 5^2

OpenStudy (anonymous):

?

OpenStudy (asnaseer):

yes 3375 you should now be able to work out what 'k' should be equal to

OpenStudy (anonymous):

oh i did its 225

OpenStudy (asnaseer):

perfect! well done!

OpenStudy (anonymous):

so basically cube all the factors and times them?

OpenStudy (asnaseer):

correct :)

OpenStudy (anonymous):

yay ty

OpenStudy (asnaseer):

yw :)